# Never Miss A Beat: An Efficient Recipe for Context Window Extension of Large Language Models with Consistent “Middle” Enhancement

### v1

@@ 3.5 Ablation Study
We compare our choice of fixing the head and tail segments with three alternatives: (i) removing both the head and tail segment, (ii) fixing only the head segment, and (iii) fixing only the tail segment (see Figure 5(b)). We find that: removing the head and tail segments leads to the worst performance; it results in a complete failure (i.e., zero score) for the context size 32K. Keeping either of the head and tail segments performs slightly better than removing both but underperforms our default choice of fixing both. We suppose that this is because fixing both gives rise to better relativity information, a finding that is consistent with that of Han et al. [23].
@@ 4 Related Works
FoT [12] addresses the limitations of local attention in transformers by integrating memory attention layers, which enable large models to learn from a wide context while reducing interference. Infini-attention [13] incorporates compressed memory into the standard attention mechanism and integrates masked local attention and long-term linear attention mechanisms within a single Transformer block. LLoCO [14] employs LoRA in conjunction with context compression, retrieval, and parameter-efficient fine-tuning to learn context offline. Although these methods can successfully extend the long context window of LLMs, they either require modifications to the attention mechanism or the addition of extra modules for assistance. In contrast, CREAM does not require these operations and can be directly applied to pre-trained model.
@@
Chen et al. [7] first proposed extending the context window through positional interpolation, which linearly reduces the input position indices to match the original context window size, thereby preventing catastrophic high attention scores from completely disrupting the self-attention mechanism. Subsequently, various methods (such as NTK [8], ABF [10], and EABF [11]) emerged that modify the base frequency of rotary positional encoding to achieve positional interpolation. YaRN [9] introduced a segmented interpolation method, applying different positional interpolations to different dimensions. LongRoPE [37] identifies and utilizes two forms of non-uniformity in positional interpolation through search, and introduces a progressive expansion strategy for positiona interpolation. Moreover, CREAM can be combined with any positional interpolation method.
@@
RandPos [17] first modified position indices so that the model leverages the relativity of positions, enabling it to extend to the target length with fine-tuning over shorter lengths. PoSE [1] then emphasized the importance of continuous segments, dividing the training length into two parts to further enhance the interpolation effect. CREAM utilizes both relativity and continuity, and it also better enables the model to focus on the middle part of the context.
@@ Appendix A Relative Positional Encoding in RoPE
We provide a simple background proof on the relative positional encoding performed by Rotary Position Embedding (RoPE) [20]. Given two embedding vectors $\bm{x}_{q},\bm{x}_{k}\in\mathbb{R}^{d}$ corresponds to query and key at positions $(m,n)\in[0,L)$, where $d$ is embedding dimension, their encoding counterparts can be defined as:
@@ Appendix C Experimental Details
We fine-tune all models by optimizing the causal language modeling objective. A learning rate of $2\times 10^{-5}$ with a linear scheduler is adopted, incorporating 10 warm-up steps. We use the AdamW [38] optimizer with the hyperparameter configurations specified by PyTorch [39]. To speed up fine-tuning, we resort to DeepSpeed  [^4] https://github.com/microsoft/DeepSpeed ZeRO stage 1 and Flash Attention-2 [40]. We perform fine-tuning on two A100-80G GPUs with a total batch size of 32 and run inference on a single A100-80G GPU. For CREAM-Base, we fine-tune it for 1,000 steps on a dataset derived from Pile [41]; for CREAM-Chat, we fine-tune it for 100 steps on ShareGPT [42]. To ensure fair comparison, we follow the fine-tuning and inference configurations established by Zhu et al. [1].
@@
For training the Base model, we directly utilize The Pile data provided by Zhu et al. [1], and select samples with token lengths exceeding 4K. For training the Chat model, we filter the ShareGPT data from public datasets [^5] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered. Specifically, we used the Vicuna prompt template to sequentially concatenate the ShareGPT data until each data point comprises at least 4K tokens. Then, we select 3.2K data points to train for 100 steps. Particularly, during the instruction tuning process, we mask the USER part and allow the model to calculate the loss only on the ASSISTANT part. We utilize two A100-80G machines with a global batch size of 32, fully utilizing the available memory. Running 1,000 steps for the Base model takes approximately 6 hours, while running 100 steps for the Chat model takes approximately 2 hours.
Appendix D LongChat Lines Results
The interpolation methods using NTK and Yarn are presented in Figures 6 and 7. As can be seen, CREAM performs the same as the Linear method for interpolation, still outperforming other methods. The result of NTK at 26K-32K is zero, which is due to the inherent properties of NTK, a finding that is aligns with Zhu et al. [1].
@@ Appendix D LongChat Lines Results
Appendix E LongBench Subtasks Results
The results of each subtask in Table LABEL:tab:longbench2 are shown in Tables  LABEL:tab:subtask1 and LABEL:tab:subtask2.
Table 6: Experimental results (%) of the LongBench subtasks selected in Zhang et al. [11]. indicates results quoted from Zhang et al. [11]. Len represents the context length during fine-tuning. All results are based on Llama 2-7B.
@@ Appendix E LongBench Subtasks Results
It is noteworthy that, to provide further evidence of the efficacy of our model, we have specifically chosen 12 tasks from the four categories outlined in Zhang et al. [11] for comparison purposes. As delineated in Table LABEL:tab:longbench1, we are able to attain superior performance on LongBench in comparison to EABF [11], even with shorter training lengths and less data.
Appendix F Limitations
When extending the context beyond the pre-trained length, there is an inevitable loss of information due to position interpolation, particularly when fine-tuning is restricted to the pre-trained length. However, in comparison to previous methods such as RandPos [17] and PoSE [1], CREAM has effectively mitigated the issue of “Lost-in-the-Middle” by introducing truncated Gaussian sampling. Additionally, as discussed in reference  Liu et al. [15], decoder-only models are prone to inherently exhibiting a U-shaped performance curve on this task. Therefore, completely solving this problem remains challenging.
Appendix G Loss Curve

### v2

@@ 3.6 Ablation Study
We compare our choice of fixing the head and tail segments with three alternatives: (i) removing both the head and tail segment, (ii) fixing only the head segment, and (iii) fixing only the tail segment (see Figure 5(b)). We find that: removing the head and tail segments leads to the worst performance; it results in a complete failure (i.e., zero score) for the context size 32K. Keeping either head or tail segments performs slightly better than removing both but underperforms our default choice of fixing both. We suppose that this is because fixing both gives rise to better relativity information, a finding that is consistent with that of Han et al. (2023).
@@
Ablation of Hyperparameters
In our implementation of truncated Gaussian sampling, as illustrated in Equation 3, the only hyperparameters are the mean $\mu$ and the variance $\sigma$. The mean $\mu$ is determined by the expansion factor. The variance $\sigma$ is adaptable based on data, we conducted experiments with five different values of $\sigma$. The results, as presented in Figure 6, indicate that the current selection ($\sigma=3$) yields optimal performance.
Figure 6: Ablation Results (%) on LongChat-Lines. Each length consisting of 50 samples. The above are the results of using Linear interpolation on the Llama 2-7B model.
@@ 4 Related Works
FoT (Tworkowski et al., 2024) addresses the limitations of local attention in transformers by integrating memory attention layers, which enable large models to learn from a wide context while reducing interference. Infini-attention (Munkhdalai et al., 2024) incorporates compressed memory into the standard attention mechanism and integrates masked local attention and long-term linear attention mechanisms within a single Transformer block. LLoCO (Tan et al., 2024) employs LoRA in conjunction with context compression, retrieval, and parameter-efficient fine-tuning to learn context offline. Although these methods can successfully extend the long context window of LLMs, they either require modifications to the attention mechanism or the addition of extra modules for assistance. In contrast, CREAM does not require these operations and can be directly applied to a pre-trained model.
@@
Chen et al. (2023) first proposed extending the context window through positional interpolation, which linearly reduces the input position indices to match the original context window size, thereby preventing catastrophic high attention scores from completely disrupting the self-attention mechanism. Subsequently, various methods (such as NTK (Peng and Quesnelle, 2023), ABF (Xiong et al., 2023), and EABF (Zhang et al., 2024)) emerged that modify the base frequency of rotary positional encoding to achieve positional interpolation. YaRN (Peng et al., 2023) introduced a segmented interpolation method, applying different positional interpolations to different dimensions. LongRoPE (Ding et al., 2024) identifies and utilizes two forms of non-uniformity in positional interpolation through search, and introduces a progressive expansion strategy for positiona interpolation. Moreover, CREAM can be combined with any positional interpolation method.
@@
RandPos (Ruoss et al., 2023) first modified position indices so that the model leverages the relativity of positions, enabling it to extend to the target length with fine-tuning over shorter lengths. PoSE (Zhu et al., 2023) then emphasized the importance of continuous segments, dividing the training length into two parts to further enhance the interpolation effect. CREAM utilizes both relativity and continuity, and it also better enables the model to focus on the middle part of the context.
@@ 5 Conclusion
Acknowledgement
The authors thank the reviewers for their insightful suggestions to improve the manuscript. This work presented herein is supported by the National Natural Science Foundation of China (62376031).
@@ Appendix A Relative Positional Encoding in RoPE
We provide a simple background proof on the relative positional encoding performed by Rotary Position Embedding (RoPE) Su et al. (2024). Given two embedding vectors $\bm{x}_{q},\bm{x}_{k}\in\mathbb{R}^{d}$ corresponds to query and key at positions $(m,n)\in[0,L)$, where $d$ is embedding dimension, their encoding counterparts can be defined as:
@@ Appendix C Experimental Details
We fine-tune all models by optimizing the causal language modeling objective. A learning rate of $2\times 10^{-5}$ with a linear scheduler is adopted, incorporating 10 warm-up steps. We use the AdamW Loshchilov and Hutter (2018) optimizer with the hyperparameter configurations specified by PyTorch Paszke et al. (2019). To speed up fine-tuning, we resort to DeepSpeed  [^4] https://github.com/microsoft/DeepSpeed ZeRO stage 1 and Flash Attention-2 Dao (2023). We perform fine-tuning on two A100-80G GPUs with a total batch size of 32 and run inference on a single A100-80G GPU. For CREAM-Base, we fine-tune it for 1,000 steps on a dataset derived from Pile Gao et al. (2020); for CREAM-Chat, we fine-tune it for 100 steps on ShareGPT Zheng et al. (2024). To ensure fair comparison, we follow the fine-tuning and inference configurations established by Zhu et al. (2023).
@@
For training the Base model, we directly utilize The Pile data provided by Zhu et al. (2023), and select samples with token lengths exceeding 4K. For training the Chat model, we filter the ShareGPT data from public datasets [^5] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered. Specifically, we used the Vicuna prompt template to sequentially concatenate the ShareGPT data until each data point comprises at least 4K tokens. Then, we select 3.2K data points to train for 100 steps. Particularly, during the instruction tuning process, we mask the USER part and allow the model to calculate the loss only on the ASSISTANT part. We utilize two A100-80G machines with a global batch size of 32, fully utilizing the available memory. Running 1,000 steps for the Base model takes approximately 6 hours, while running 100 steps for the Chat model takes approximately 2 hours.
Appendix D Robustness Across LLMs
Our proposed method exhibits strong generalization capabilities and can be applied to other large language models (LLMs) without the need for parameter modification. To validate this, we conducted experiments on Baichuan2-7B, with the corresponding results presented in Table LABEL:tab:baichuan2.
Table 7: Perplexity results of GovReport and Proof-pile. Each experiment is the average perplexity of 50 samples, and all results are based on Baichuan2-7B fine-tuned on 4K data length.
Model GovReport Proof-pile
4K 8K 16K 32K 4K 8K 16K 32K
Original 3.3 - - - 5.8 - - -
CREAM-Linear 3.6 2.9 2.5 2.2 6.2 6.1 6.0 5.8
Furthermore, we fine-tuned LLaMa3-8B using a context window size of $4K$ tokens, with the experimental outcomes shown in Table LABEL:tab:llama3.
Table 8: Results (%) on LongChat-Lines. Each length consists of 50 samples. All results are fine-tuned on Llama-3-8B with 4K length data through linear position interpolation.
AVG Length 2000 4000 7800 8800 9700 11000 12000 14000 17000 19000 24000 28000 32000
CREAM-Linear 0.98 1.00 0.96 0.94 0.86 0.92 0.92 0.92 0.86 0.84 0.70 0.60 0.48
The results in Tables LABEL:tab:baichuan2 and LABEL:tab:llama3 clearly demonstrate the transferability of our method to different models, underscoring its robustness. Of particular note is that despite LLaMa3-8B having a native context length of $8K$ tokens, fine-tuning on training data with a $4K$ context window yielded unexpectedly strong performance.
Appendix E LongChat Lines Results
The interpolation methods using NTK and Yarn are presented in Figures 7 and 8. As can be seen, CREAM performs the same as the Linear method for interpolation, still outperforming other methods. The result of NTK at 26K-32K is zero, which is due to the inherent properties of NTK, a finding that is aligns with Zhu et al. (2023).
@@ Appendix E LongChat Lines Results
Appendix F LongBench Subtasks Results
The results of each subtask in Tables LABEL:tab:longbench2 are shown in Tables  LABEL:tab:subtask1 and LABEL:tab:subtask2.
Table 9: Experimental results (%) of the LongBench subtasks selected in Zhang et al. (2024). indicates results quoted from Zhang et al. (2024). Len represents the context length during fine-tuning. All results are based on Llama 2-7B.
@@ Appendix F LongBench Subtasks Results
It is noteworthy that, to provide further evidence of the efficacy of our model, we have specifically chosen 12 tasks from the four categories outlined in Zhang et al. (2024) for comparison purposes. As delineated in Table LABEL:tab:longbench1, we are able to attain superior performance on LongBench in comparison to EABF Zhang et al. (2024), even with shorter training lengths and less data.
Appendix G Limitations
When extending the context beyond the pre-trained length, there is an inevitable loss of information due to position interpolation, particularly when fine-tuning is restricted to the pre-trained length. However, in comparison to previous methods such as RandPos Ruoss et al. (2023) and PoSE Zhu et al. (2023), CREAM has effectively mitigated the issue of “Lost-in-the-Middle” by introducing truncated Gaussian sampling. Additionally, as discussed in reference  Liu et al. (2024), decoder-only models are prone to inherently exhibiting a U-shaped performance curve on this task. Therefore, completely solving this problem remains challenging.
Appendix H Loss Curve